# Data Analysis Prediction over Multiple Unseen Datasets: A Vector Embedding Approach

### v2

@@ Abstract
The rapid growth in data volume and the increasing availability of datasets present analysts with the challenge of selecting high-quality data to optimize the performance of analytics operations. While choosing the best-quality data significantly improves accuracy and efficiency, it remains a complex task, particularly when dealing with a vast number of inputs. To tackle this challenge, we introduce a novel approach that predicts analytics outcomes by building a model based on datasets similar to the target dataset. Our method determines dataset similarity by mapping each dataset into a vector embedding representation. This transformation is achieved through our deep learning model, NumTabData2Vec, which processes entire datasets and projects them into a lower-dimensional vector space. Through experimental evaluation, we compare our framework’s prediction accuracy and execution time against a leading state-of-the-art modelling operator, demonstrating that our approach effectively predicts analytics outcomes. Additionally, our vectorization model can capture and differentiate diverse real-world scenarios by representing them in a compact vector space.
Index Terms:
Analytics Modelling, Vector embeddings, Vector Similarity, Data Quality
I Introduction
@@
Big data technologies daily face the rapid evolution in volume as well as variety and velocity of processed data [1]. Such big data characteristics routinely force analytics pipelines to underperform, requiring continuous maintenance and optimization. One major reason for this is bad quality of data [^1] https://tinyurl.com/de62sf48. Poor data quality leads to low data utilisation efficiency and even brings forth serious decision-making errors[2].
@@
Data quality can be improved when focusing on the actual content of the data. Data-centric Artificial Intelligence (AI) [3] emphasises on the quality, context, and structure of the data to improve its quality, as well as the analytical or machine learning (ML) algorithmic performance. Understating the data context properties such as data features, origins, relevance, and potential biases plays a critical role in modelling more accurate and reliable models. Data-centric AI prioritises the process of refining and enriching datasets to make them more suitable for real-world applications. In a similar way, a lot of scientists make the case that improving data quality by focusing on the content is a crucial factor in achieving better results [3].
@@
A plethora of available data sources and datasets in the organisations’ data centres and lakes pose a significant challenge: Deciding which datasets should be selected and passed into analytic workflows in order to produce more accurate results/predictions. While modern analytics workflows have evolved to incorporate diverse operators, the process of selecting the most suitable datasets using data-centric AI techniques to optimise a possibly complex computation is still under research [4]. When dataset selection is left to human experts, it can result in low prediction performance and can be time-consuming. Equally costly can be the process of exhaustively executing over all (or many) available dataset inputs, in order to identify high-quality inputs according to business standards.
@@
In previous work [12], predicting the output of an analytics operator assuming a plethora of available input datasets was tackled via the creation of an all-pair similarity matrix, which, relative to the similarity function used, reflected the distance between datasets over a single characteristic (e.g., data distribution).
To improve the big data analysis and utilize modern AI methods and systems, data or vector embeddings have been suggested. Data embedding vectorization [14, 15] aims at projecting data from a high-dimensional representation space into a more compact, lower dimensional space. By extracting meaningful patterns from the data features using deep learning, data is projected to a lower representation space. Using this space, one can find the most similar data points among a dataset.
@@
For the purpose of improving the accuracy of a modelled analytic operator (i.e., predict the outcome of a ML algorithm without actually executing it due to its cost), we propose a modelling framework that employs vector embeddings for dataset selection from a data lake store that contains a large number of available datasets. In order to predict an operator’s output for an “unseen” query dataset, our method considers only datasets that are qualitatively similar to the query, utilising a similarity search over the vector embeddings. The selection of similar datasets reduces the prediction error, as well as the cost to model the operator, under the assumption that realistic analytical operators perform similarly under similar inputs. Similarity search is thus executed on the dataset vector embedding representations in a low-dimensional space. Each vector embedding is generated using our proposed deep learning method, NumTabData2Vec, which takes an entire dataset as input and transforms it into a vector embedding representation space. While previous methods split the dataset into chunks and produced a vector for each chunk or used the dataset metadata information, we use the entire dataset utilising all data record values. Our framework thus models an operator using similar datasets to the queried one. An interesting feature of our methodology is the ability of the NumTabData2Vec to efficiently distinguish with detail between different datasets and record properties in the lower vector representation space. Furthermore, the same vector representation is used for the modelling of multiple analytics operators, offering enhanced flexibility and time-efficiency. Compared to similar previous work [12, 13], our work uses state-of-the-art data representation (vector embeddings) which are able to capture multiple data properties that can be used in order to assess similarity, namely record order, dataset size, data distribution, etc.
@@
  • We introduce a framework for operator modelling in order to predict its outcome on an unseen tabular input dataset from a plethora of available ones. Our method uses dataset vector embedding representations to improve the prediction performance via selecting the most relevant datasets to base its prediction upon.
@@
  • We provide an experimental evaluation of our proposed methodology using multiple real-world scenarios and compare it directly to system Apollo [12, 13].
Our evaluation illustrates that our proposed methodology can model operators with low prediction error by adaptively selecting similar quality datasets for each model. Moreover, it is able to do so with considerable amortized speed-ups. The evaluation of NumTabData2Vec indicates that it accurately projects any available dataset into a vector embedding representation and accurately models different dataset properties within the representation space.
II Related Work
Prior efforts largely concentrate on boosting/scaling an algorithm’s performance utilizing only the size of data input (record number) without examining its quality. As such, we also discuss works that try to identify the most suitable data features for the optimisation of analytic task operators. Vectorising data to lower embedding representation is a modern method that helps in identifying significant features across data types and datasets. As vector embeddings extract important features from data, we discuss studies that used the feature representation of data tuples to improve ML model prediction.
II-A Data Quality
Big data applications aim to improve data quality by addressing various challenges. Dagger [8] enhances data quality by detecting errors in data pipelines using an SQL-like language and allowing code-handling for variables and arrays. ReClean [9] automates tabular data cleaning via reinforcement learning, optimizing tool selection for error repair. IterClean [10] employs a large language model (LLM) to iteratively clean data by labelling initial tuples and using error detection, verification, and repair. In [11], data tuple quality is measured using Shapley values from game theory, with Truncated Monte Carlo Shapley and Gradient Shapley methods estimating a tuple’s value to a learning algorithm. Apollo [12, 13] is a content-based method that predicts the outcome of an analytic operator by leveraging dataset similarity. It involves three steps: creating a similarity matrix, projecting datasets to a lower-dimensional space, and modelling the operator using a small random subset of datasets. Unlike Apollo, our approach selects the most relevant, high-quality datasets to model analytic operators, aiming to improve prediction performance. Moreover, modelling and similarity occur over the vector embeddings which are able to incorporate all dataset properties. In contrast, similarity functions used in [12] can only target a single property.
II-B Dataset Selection Inference
@@
II-C Data Vectorization and Embeddings
The goal of data vectorization is to project high-dimensional data into a lower-dimensional vector space. Word2Vec [14] generates word embeddings using two neural network architectures, Continuous Bag of Words and Skip-Gram, leveraging word context. Graph2Vec [15] creates graph embeddings by dividing graphs into sub-graphs with a skip-graph model and aggregating their embeddings. ImageDataset2Vec [17] extracts meta-features from image datasets to generate embeddings, helping to select the most suitable classification algorithm. Dataset2Vec [18] uses meta-features and the DeepSet model to project datasets into embeddings and measure dataset similarity. Table2Vec [19] generates table embeddings by incorporating data features, metadata, and structural elements like captions and column headings. Inspired by these methods, we designed a model that generates vector representations of tabular datasets over their record data values, not their metadata.
Vector embeddings, which capture valuable information from data tuples, are widely used in classification tasks. TransTab [20] transforms data tuples into vector embeddings, applying transformer layers to encode features and predict classes, enhanced by supervised and self-supervised pretraining. FT-Transformer [21] and Res-Net architectures similarly use embeddings of categorical and numerical features, processed through transformer layers for class prediction. Tab-Transformer [22] embeds categorical features into vectors, normalizes continuous features, and combines them in a Multi-Layer Perceptron (MLP) to predict classes. Unlike these methods, which focus on tuple-level embeddings, our framework uses dataset-level vector embeddings to identify the most relevant datasets, improving analytic operator performance.
III Methodology
@@
III-A Framework Architecture
Consider a data lake store that contains a (possibly large) number $n$ of datasets $D=\left(d_{1},\ d_{2},\ d_{3},\ \ldots,\ d_{n}\right)$. Also, let us consider an analytics operator (e.g., a ML algorithm) $\Phi$ and an “unseen” dataset $D_{o}$. We assume that each $D_{i},1\leq i\leq n$ as well as $D_{o}$ consist of records with the same number of columns with numerical values only. Each dataset can, naturally, have a different number of records. Operator $\Phi$ consumes a single such dataset as input to produce a single numerical output: $\Phi:D_{i}\rightarrow\mathbb{R}$. Our goal is to predict $\Phi(D_{o}),\forall D_{o}$ with a minimal cost and prediction error. Our premise is that the output of an analytical operator on any unseen $D_{o}$ can be accurately modelled and runtime efficiency improved by selecting a small subset of $D$. This can be achieved by identifying the most similar datasets to $D$, denoted: $D_{o}$: $D_{r}\subseteq D_{o}$.
@@ III-A Framework Architecture
Figure 1 depicts the pipeline of our proposed framework. Datasets in $D$ are transformed into $k$-dimensional vectors via our NumTabData2Vec embedding scheme. Dataset vectorization needs no recomputation as the vectors are saved and can be reused for the rest of the framework steps. Each time a new dataset $D_{o}$ needs to be inferred relative to an analytics operator $\Phi$, its vector embedding is equally created. The datasets used for the creation of the model are selected via similarity search. Using different similarity functions, a small subset of datasets similar to $D_{o}$ are identified and used in order to create a model that predicts $\Phi(D_{o})$. With this approach, our framework is utilizing “right quality” data in its inference mechanism, with irrelevant and extraneous datasets being excluded from the modelling process.
The datasets are embedded by our NumTabData2Vec method. It takes an entire dataset as input and transforms it into a $k$-dimensional vector embedding representation space named $z$. The Vector embedding $z$, is a lower dimension representation of the dataset with the entire characteristics of the dataset being represented. Each dataset from $D$ is projected into such a $k$-dimensional vector embedding representation. The datasets are consequently represented by $D_{z}$ which contains all vector representations $\left(z_{1},\ z_{2},\ z_{3},\ \ldots,\ z_{n}\right)$. Dataset $D_{o}$ is also vectorized into $z_{o}$. Using the embedding representation $z_{o}$ and applying different similarity functions over the vector representations $D_{z}$, we may choose the most similar subset of $D$. The final step of the pipeline involves the operator modelling with any relevant method (e.g., Linear Regression, SVM, Multi-Layer Perceptron, etc.). This model can be used in order to infer the value of $\Phi(D_{o})$.
III-B NumTabData2Vec
@@
This method transforms a dataset $D_{i}$ to a lower dimension representation $z$ by using only the numerical values of the dataset’s records and excluding metadata, column names, and filenames. We present a deep learning model architecture based on the variational autoencoder (VAE) [5] concept. For each dataset, data is projected from a high dimensional space to a lower vector embedding representation space $z$, with $1\times k$ dimension. The proposed method is thus defined as:
$NumTabData2Vec\left(D_{i}\right)\rightarrow z,$ (1)
where the model takes an $m\times n$ dimensional numerical dataset and projects it to a lower $k$-dimension space $z$ ($k>1$). We wish for our method to be generally applicable and able to use any scenario dataset that learns to project a vector embedding representation during the training phase. Also, we wish this method to be able to learn this vector embedding representation using a diverse amount of data during the training phase, and execute thereafter without the need of further training, fine-tuning, or online-training methods. Finally, we need our scheme to be able to quickly and precisely extract the vector for every dataset input as well as deal with varying dataset dimensions without requiring any modifications.
The deep learning model architecture is depicted in Figure 2, where a dataset $D_{i}$ with dimensions $m\times n$ passes through the encoder and is projected into a vector embedding representation $z$. The dataset has to be recreated since the decoder uses the same layers as the encoder but in the opposite direction. The learning of the vector representation $z$ is using the probabilistic encoder $q_{\phi}\left(z|x\right)$ distribution and the probabilistic decoder $p_{\theta}\left(x|z\right)$ distribution utilising the Kullback–Leibler (KL) divergence [6]. To achieve that, the following condition:
$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)$ (2)
of Kullback–Leibler (KL) divergence must be minimised. To learn the new vector representation $z$, the dataset must be reconstructed back from $z$ to their input format using the decoder, to verify if the vector representation is on a lower dimension in terms of the input dataset. This is the reconstruction loss. The loss function can be described as:
@@
$argmax\mathcal{L}_{\theta,\phi}\left(x\right)$ (4)
@@
The decoder takes as an input the dataset $D_{i}$, first extracts the feature embeddings from the input layer and passes them into the Transformer layers. For the transformer layer we are using the pre-LN Transformer layer [7] instead of the traditional post-LN transformer layer where the normalisation layer is employed inside the residual connection and before the prediction of the feed-forward layer. Our model design takes into account and uses the Pre-LN Transformer layer as it demonstrates improved learning outcomes with shorter training times [7]. Following that, the transformed embedding space is projected onto a probabilistic distribution vector space $z$ using the mean ($\mu$) and standard deviation ($\sigma$). Then the transformed embedding space using $\mu$ and $\sigma$ are projected into a probabilistic distribution vector space $z$. The vector embedding space $z$ is a lower-dimensional representation of the dataset $D_{i}$ that preserves all the essential information needed to describe the dataset within the space $z$. The higher the $k$-dimension of the $z$ vector embedding representation, the higher and more accurate is this representation, as more features are extracted from the dataset [18].
III-C Dataset Selection

### v3

# Analytics Modelling over Multiple Datasets using Vector Embeddings

@@ Abstract
The massive increase in the data volume and dataset availability for analysts compels researchers to focus on data content and select high-quality datasets to enhance the performance of analytics operators. While selecting high-quality data significantly boosts analytical accuracy and efficiency, the exact process is very challenging given large-scale dataset availability. To address this issue, we propose a novel methodology that infers the outcome of analytics operators by creating a model from the available datasets. Each dataset is transformed to a vector embedding representation generated by our proposed deep learning model NumTabData2Vec, where similarity search are employed. Through experimental evaluation, we compare the prediction performance and the execution time of our framework to another state-of-the-art modelling operator framework, illustrating that our approach predicts analytics outcomes accurately, and increases speedup. Furthermore, our vectorization model can project different real-world scenarios to a lower vector embedding representation accurately and distinguish them.
Keywords:
Data Quality Analytics Modelling Vector embeddings Vector Similarity
1 Introduction
@@
Big data technologies daily face the rapid evolution in volume as well as variety and velocity of processed data [12]. Such big data characteristics routinely force analytics pipelines to underperform, requiring continuous maintenance and optimization. One major reason for this is bad data quality [^1] https://tinyurl.com/de62sf48. Poor data quality leads to low data utilisation efficiency and even brings forth serious decision-making errors[6].
@@
Data quality can be improved when focusing on the actual content of the data. Data-centric Artificial Intelligence (AI) [30] emphasises on the quality, context, and structure of the data to improve its quality, as well as the analytical or machine learning (ML) algorithmic performance. Understanding the data context properties, such as data features, origins, relevance, and potential biases, plays a critical role in modelling more accurate and reliable models. Data-centric AI prioritises the process of refining and enriching datasets to make them more suitable for real-world applications. Similarly, many researchers argue that prioritizing content-focused data quality is essential for achieving superior results [30].
@@
Yet, the plethora of available data sources and datasets in an organisation data repository poses a significant challenge: Deciding the most suitable datasets for analytics workflows to ensure accurate results/predictions. While modern analytics workflows incorporate diverse operators, optimising dataset selection using data-centric AI methods remains an active research area [15]. When dataset selection is left to human experts, prediction performance drops, and it consumes more time. Equally costly and inefficient is the evaluation of all available datasets to identify high-quality inputs.
@@
In previous work [9], predicting the output of an analytics operator assuming a plethora of available input datasets was tackled via the creation of an all-pair similarity matrix, which, relative to the similarity function used, reflected the distance between datasets over a single data quality metric (e.g., data distribution). Data or vector embeddings have been proposed to enhance big data analysis and modern AI systems. Data embedding vectorization [22, 24] aims at projecting data from a high-dimensional representation space into a more compact, lower-dimensional space. Extracting meaningful information through data features using deep learning, data is projected to a lower representation space.
@@
To improve the accuracy of a modelled analytic operator (i.e., predict the outcome of a ML algorithm without actually executing it due to its cost), we propose a framework that uses vector embeddings for dataset selection from a large data lake repository. Our method predicts an operator’s output for an “unseen” query dataset, by selecting qualitatively similar datasets through similarity search over the vector embeddings. The selection of similar datasets reduces the prediction error, as well as the cost to model the operator, under the assumption that realistic analytical operators perform similarly under similar inputs. The embeddings are generated using our deep learning method, NumTabData2Vec, which processes entire tabular datasets rather than chunks or metadata, enabling efficient distinction between datasets and flexible modelling of multiple operators. Compared to similar previous work [9, 4], our work uses state-of-the-art data representation (vector embeddings) which are able to capture multiple data properties that can be used in order to assess similarity, namely record order, dataset size, data distribution, etc.
@@
  • We introduce a framework for operator modelling (open-source prototype [^2] GitHub Repository) in order to predict its outcome on an unseen tabular input dataset from a plethora of available ones. Our method uses dataset vector embedding representations to improve the prediction performance via selecting the most relevant datasets to base its prediction upon.
@@
  • We provide an experimental evaluation of our proposed methodology using multiple real-world scenarios and compare it directly to the Apollo system [9, 4].
Our evaluation illustrates that our methodology produces low prediction error by adaptively selecting similar quality datasets, achieving significant amortized speed-ups. NumTabData2Vec evaluation shows that it effectively projects datasets into vector embeddings while accurately capturing diverse dataset properties within the representation space.
2 Related Work
Prior efforts have focused on boosting algorithm performance by increasing data input (record number) rather than assessing quality. Consequently, we review works that identify optimal data features for analytic operator optimization. Vectorising data to lower embedding representation is a modern method that helps in identifying significant features across data types and datasets. As vector embeddings extract important features from data, we discuss studies that used the feature representation of data tuples to improve ML model prediction.
2.1 Data Quality
Big data applications aim to improve data quality by addressing various challenges. Dagger [26] enhances data quality by detecting pipeline errors using an SQL-like language, while ReClean [3] automates tabular data cleaning via reinforcement learning. IterClean [25] employs a large language model (LLM) to iteratively clean data by labelling initial tuples and using error detection, verification, and repair. In [8], data tuple quality is measured using Shapley values from game theory, with Truncated Monte Carlo Shapley and Gradient Shapley methods estimating a tuple’s value to a learning algorithm. Apollo [9, 4] is a content-based method predicts analytic operator outcomes by leveraging dataset similarity through three steps: creating a similarity matrix, projecting datasets to a lower-dimensional space, and modelling the operator using a small random subset of datasets. Unlike Apollo, our approach selects the most relevant, high-quality datasets to model analytic operators, aiming to improve prediction performance. Additionally, our vector embeddings incorporate all dataset properties, whereas Apollo’s [9] similarity functions target only a single property.
2.2 Dataset Selection Inference
@@
2.3 Data Vectorization and Embeddings
The goal of data vectorization is to project high-dimensional data into a lower-dimensional vector space. Word2Vec [22] (using Continuous Bag of Words and Skip-Gram) [22] leverages word context to generate embeddings. Graph2Vec [24] creates graph embeddings by dividing graphs into sub-graphs with a skipgram model and aggregating their embeddings. ImageDataset2Vec [7] extracts meta-features from image datasets to generate embeddings, helping to select the most suitable classification algorithm. Dataset2Vec [16] uses meta-features and the DeepSet model to project datasets into embeddings and measure dataset similarity. Table2Vec [31] generates table embeddings by incorporating data features, metadata, and structural elements like captions and column headings. Mix2Vec [32], is an unsupervised deep neural network that projects mixed data into vector embeddings. In a clustering experiment like in their work on the common Adult dataset, our model outperformed Mix2Vec (recent method without publicly available code) by nearly $10\%$, demonstrating superior performance. Inspired by these methods, we designed a model that generates vector representations of tabular datasets over their record data values, not their metadata.
Vector embeddings, which capture valuable information from data tuples, are widely used in classification tasks. TransTab [28] encodes features with transformer layers to predict classes, leveraging supervised and self-supervised pretraining. FT-Transformer [10] and Res-Net architectures similarly use embeddings of categorical and numerical features, processed through transformer layers for class prediction, while Tab-Transformer [14] combines embedded categorical and normalized continuous features in an MLP for class prediction. Unlike these tuple-level approaches, our framework uses dataset-level embeddings to identify relevant datasets, enhancing analytic operator performance.
3 Methodology
@@
3.1 Framework Architecture
Consider a data lake repository that contains a (possibly large) number $n$ of structured tabular datasets $D=\left(d_{1},\ d_{2},\ d_{3},\ \ldots,\ d_{n}\right)$. Also, let us consider an analytics operator (e.g., a ML algorithm) $\Phi$ and an “unseen” dataset $D_{o}$ (from the same domain). We assume that each $D_{i},1\leq i\leq n$ as well as $D_{o}$ consist of records with numerical values only. Each dataset can, naturally, consist of different number of records. Operator $\Phi$ consumes a single such dataset as input to produce a single numerical output: $\Phi:D_{i}\rightarrow\mathbb{R}$. Our goal is to predict $\Phi(D_{o})$ with minimal cost and error by modelling the operator’s output for $D_{o}$ using a small subset of similar datasets $D_{r}\subseteq D$.
@@ 3.1 Framework Architecture
Figure 1 depicts the pipeline of our proposed framework. Datasets in $D$ are transformed into $k$-dimensional vectors using our NumTabData2Vec scheme, and these embeddings are stored for reuse. Each time a $D_{o}$ needs to be inferred relative to an analytics operator $\Phi$, its vector embedding is created. The datasets used for the creation of the model are selected via similarity search to produce a small subset of relevant datasets. These chosen datasets are then used to model and predict $\Phi(D_{o})$, ensuring that only high-quality, pertinent data is processed. With this approach, our framework is utilizing “right quality” data in its inference mechanism, with irrelevant and extraneous datasets being excluded from the modelling process.
The datasets are embedded by our NumTabData2Vec method, which transforms each entire dataset in $D$ into a k-dimensional vector $z$ that captures all its characteristics. Our framework operates seamlessly across diverse real-world scenarios without modification, requiring only the specification of a repository containing distinct numerical tabular datasets. The Vector embedding $z$ is a lower-dimension representation of the dataset with the entire characteristics of the dataset being encoded. Dataset $D_{o}$ is similarly embedded as $z_{o}$. Using the embedding representation $z_{o}$ and applying different similarity functions over the vector representations $D_{z}$, we may choose the most similar subset of $D$. The final step of the pipeline involves the operator modelling with any relevant method (e.g., Linear Regression, SVM, Multi-Layer Perceptron, etc.). This model is then used in order to infer the value of $\Phi(D_{o})$.
@@
3.2 NumTabData2Vec
This method transforms a dataset $D_{i}$ into a lower-dimensional vector $z$ using only its numerical values while excluding metadata like column names and filenames. We present a deep learning model architecture based on the variational autoencoder (VAE) [17] concept, that projects high-dimensional data into a $1\times k$ vector embedding dimension. The proposed method is thus defined as:
$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\$ (1)
where the model takes an $m\times n$ dimensional numerical dataset and projects it to a lower $k$-dimension space $z$ ($k>1$). We desire our method to be generally applicable to any dataset by learning to project a vector embedding during training. We also expect this method to learn vector embeddings from diverse data and operate without additional training or fine-tuning. Finally, we ought our scheme to be able to quickly and precisely extract vectors from every input while handling varying dataset dimensions without modifications.
The deep learning model architecture is depicted in Figure 2, where a dataset $D_{i}$ with dimensions $m\times n$ passes through the encoder and is projected into a vector embedding representation $z$, then reconstructed by a decoder that mirrors the encode. The vector representation $z$ is learned using a probabilistic encoder $q_{\phi}\left(z|x\right)$, and decoder $p_{\theta}\left(x|z\right)$ that learns the distribution utilising the Kullback–Leibler (KL) divergence [18]. To achieve that, the following condition:
$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\$ (2)
of Kullback–Leibler (KL) divergence must be minimised. To learn the new vector representation $z$, the dataset must be reconstructed back from $z$ to its input format using the decoder, to verify that the vector representation is compact. This reconstruction loss is part of the overall loss function, defined as:
@@ 3.2 NumTabData2Vec
$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\$ (4)
@@
The decoder extracts feature embeddings from the input dataset $D_{i}$ and processes them through Transformer layers. For the transformer layer we are using the pre-LN Transformer layer [29] instead of the traditional post-LN Transformer layer where the normalisation layer is employed inside the residual connection and before the prediction of the feed-forward layer. Following that, the transformed embedding space is projected into a probabilistic vector space z using the mean ($\mu$) and standard deviation ($\sigma$). This lower-dimensional space retains all essential information about $D_{i}$, and a higher dimension $k$ in $z$ leads to a more accurate representation by capturing additional features [16].
3.3 Dataset Selection